A rectangle of has an area of 48 and a perimeter of 32. Find the dimensions of the rectangle.
Answer by edjones(8007)

Lw=48
2L+2w=32
.
L+w=16
L=16-w
w(16-w)=48
16w-w^2=48
w^2-16w+48=0
(w-12)(w-4)=0
w=4, L=12
.
Ed